# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

### v4

@@ 3.1 Experiment Settings
In this study, we evaluated twelve up-to-date commercial models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), and Gemini (Gemini Team, 2024) series using our framework. Additionally, three Llama models (Touvron et al., 2023), with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and recognized capability in reasoning tasks. However, due to their limited performance on our tasks, as shown in Table 2, we focused our investigation solely on more powerful closed-source models. We included the results of open-source models for completeness.
@@
We adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024) datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks. We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
@@
MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges. The prompt templates we used are identical to those used in the benchmarks. More details about the models, tasks, and prompts can be found in Appendix. Sec. H.
@@
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials. This resulted in 576 instances per judge for MTBench and 432 instances per judge for DevBench. The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, approximately 100,000 evaluation instances were analyzed in this study.
To identify significant factors contributing to position bias, we performed bidirectional stepwise regression on data from the two benchmarks. We used variables such as average lengths of input, output, and prompt; answer quality gap; LLM judge series; candidate identities; and task categories to predict $PC$ and $PF$, respectively. Each model prunes non-significant variables based on the Akaike Information Criterion (AIC) score. This process involves both forward selection and backward elimination, with each ”step” testing whether including or excluding a variable improves the model’s AIC value. Further details about the process can be found in Appendix. Sec. F.
3.2 Empirical Results and Interpretations
@@
The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations. Fig. 3 and Fig. 7 demonstrate that position bias varies by judges and tasks significantly. Fig. 4 further investigates the impact of answer quality gap and lengths of prompt components on position bias. These analyses were conducted by considering all judges together on MTBench. The detailed analyses for each judge and on DevBench can be found in Appendix.Sec.E, and Sec. G.
Through bidirectional stepwise regression, we found that, based on data from the two benchmarks, LLM judge series, candidate identities, and task categories significantly impact Position Consistency among all variables. Similarly, these factors also contribute significantly to Preference Fairness. Additionally, we found that average output length is a statistically significant predictor of $PF$. This finding is not surprising, as longer outputs are generally perceived as higher quality and more preferred. Quantitative results and more discussions can be found in Appendix. Sec. F.
@@
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Positional Fairness ($PF$). Top 5 performances are marked in bold. “Error” is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format. High error rates and low $RS$ are marked red, implying further evaluations (i.e., $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data. Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances. In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations. Chatbot Arean Leaderboard’s overall ranking of the models is listed.
Judge MTBench DevBench Arena
@@
Caude-3.5-Sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00 7
Claude-3-Opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00 17
Claude-3-Sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00 37
Claude-3-Haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00 46
Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13 12
Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96 25
Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00 66
o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00 2
GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00 6
GPT-4-Turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00 17
GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00 52
GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00 94
Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 99
Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00 109
Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94 57
Figure 3: Baseline comparisons of judges across tasks on MTBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be GPT-4 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
@@
4 Main Findings
@@
Position Bias of Capable Judges are not Mere Random Variations As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal ”Error” rates, generally exhibit $RS$ values above 0.85. The most up-to-date and capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, GPT-4, and GPT-4o, all achieve near-perfect $RS$ scores exceeding 0.95 on both benchmarks. These results confirm that LLM judgments, and the resulting position bias, are not merely random variations. This strengthens confidence that one-time generated judgments by these validated LLMs accurately reflect their judging capabilities.
@@
Position Bias Varies by Judge & Task As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig. 3 for $PC$ comparisons. However, certain models achieve comparable or superior performances than GPT-4 on certain tasks. For example, according to Fig. 3, for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges. Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
@@
Additionally, significant variations of $PF$ across tasks are observed. The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on the others. Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, role play, and writing tasks. Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
@@
Moreover, high position consistency does not guarantee fairness. For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively. In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency.
Therefore, the position bias of LLM judges is judge-dependent and task-dependent. This observation is confirmed by the bidirectional stepwise regression where judge identities and task categories are statistically significant predictors of $PC$ and $PF$. In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance. Moreover, as shown in Table 2, a high ranking in the overall reasoning performance of a model does not guarantee its judging reliability. Similar findings can be observed in the DevBench results, as detailed in Appendix.Sec. G.
@@
Position Bias Correlates to Answer Quality Gap Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$). Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1. Fig. 4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. E.
@@
Similarly, as shown in Fig. 1(b), judgments generally become more preference fair as $\delta_{q}$ increases. However, the extent is not as significant as for $PC$. Also, the relationship varies by judge, as some LLMs maintain preference fairness regardless of $\delta_{q}$. For example, as shown in Fig. 6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
@@
These observations align with the right-arrow shape as demonstrated in Fig. 4(c), where there is a general trend that judgments become preference fairer as position consistency increases. It also justifies the reasonableness of our quantification of preference fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
@@
Our investigation also helps to explain the observed verbosity/length bias (Liu et al., 2024; Kim et al., 2023; Zheng et al., 2024b) and self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b) of LLM judges. From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities. Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the solutions generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4). In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
@@
Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length). Since the tokenization methods vary by models, we apply the length of the prompt component strings for uniformity and scalability. The analysis from Fig. 4(d)(e)(f) shows that there is very weak relationship between the lengths of prompt components and position bias. By stepwise regression, the average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table. 4. Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
@@
Practical Implications for Benchmark Design We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appendix.Sec. C to gain further insights. In particular, the disagreement analysis helps identify instances where it is easy or difficult for judges to reach a consensus, reflecting the complexity of the judgment task. Based on our findings regarding the answer quality gap and LLM agreement/disagreement, this study offers practical implications for designing evaluator benchmarks that control the difficulty of judging tasks. Our findings show that the hardest-to-judge instances are those where (1) LLMs frequently disagree with each other, (2) comparable win rates and minimal quality gaps exist between the two candidate models, and (3) significant position bias is exhibited by the majority of judges.
@@
One qualitative analysis example of a difficult-to-judge instance comes from MTBench question 120, which involved a math problem where both candidate models, Vicuna-13b-v1.3 and WizardLM-30b, provided incorrect solutions. The problem was: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find $x$ such that $f(x)=0$. Vicuna-13b-v1.3 incorrectly applied the quadratic formula to solve the cubic equation, while WizardLM-30b used an enumeration method but failed to identify the correct answer. This resulted in a 4{A}-3{B}-5{C} choice pattern among the twelve judges, illustrating a typical case of significant disagreement.
@@
We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer. Others preferred either correctness (choosing {C} as a tie) or methodological soundness (choosing {B}, favoring WizardLM-30b). In the swapped order setting, 9 out of 12 judges displayed position bias. This highlights the challenge of judging solutions with minimal quality gaps, leading to greater position bias and subjectivity.
The key takeaway is that reducing hard-to-evaluate, highly subjective instances can improve benchmark design. Likewise, minimizing trivial cases where judges easily agree would enhance the benchmark’s robustness. Clearer and more specific evaluation criteria could also help resolve ambiguity in difficult instances.
5 Conclusion
In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of LLM judges in pairwise comparative scenarios in terms of repetition stability, position consistency, and preference fairness. Through comprehensive evaluations of 12 judges on two benchmarks across 22 tasks with over 100,000 evaluation instances, we observe significant variations of position bias across judges and tasks. We also discover that position bias is weakly length-dependent but significantly impacted by the answer quality gap. Our findings enhance the understanding of position bias, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Appendix Table of Contents
[…]
[appendices]0
Appendix
Appendix A Contribution & Limitation

### v5

@@ C.1 Experiment Settings
In this study, we evaluated twelve up-to-date commercial models from the GPT [13], Claude [14], and Gemini [15] series using our systematic framework. Three Llama models [16], with parameters ranging from 7B to 13B, were selected as open-source exemplars due to their popularity and well-known capability in reasoning tasks. However, due to their limited performance on our tasks, as shown in Table 2, we focus our investigation solely on more powerful closed-source models. These models are more widely used and reliable, though we still include open-source models to gain more comprehensive insights.
@@
We adopted the modified MTBench [3] and DevBench [4] datasets for our study due to their demonstrated high human-LLM agreement and the reliability of state-of-the-art LLMs on the evaluation tasks. We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of responses.
@@
MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired these candidate models with the baseline candidate for evaluation by the LLM judges. The prompt templates we used are identical to those used in the benchmarks. More details about the models, tasks, and prompts can be found in Appendix. Sec. L.
@@
To compute repetition stability, we sampled 3 questions per task and 4 candidate models, paired with baseline candidates, for each judge to evaluate across 3 repetitive trials. This resulted in 576 and 432 instances per judge for MTBench and DevBench, respectively. The temperature hyperparameter was set to 1 for all judge models to generate nontrivial results. To compute position consistency and preference fairness, the number of instances increased to 4,800 and 2,240, covering the entire MTBench and DevBench datasets. In total, approximately 100,000 evaluation instances were analyzed in this study.
C.2 Empirical Results and Interpretations
@@
The evaluation results of 12 close-source and 3 open-source models in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not merely random variations.
This further necessitates the need for investigating and analyzing preference fairness simultaneously with position consistency. Fig. 3 and Fig. 7 demonstrate that position bias varies by judges and tasks significantly. Together with Table 2, it stresses the trade-off between $PC$ and $PF$ when selecting task-specific LLM judges. Fig. 4 further investigates the impact of answer quality gap and lengths of prompt components on position bias. These analyses are conducted by considering all judges together on MTBench. The detailed analyses for each judge and on DevBench can be accessed from Appendix.Sec.J, and Sec.K.
@@
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Positional Fairness ($PF$). Top 5 performances are marked in bold.“Error" is due to failure of judgment generation (e.g, exceeding context window) or invalid extraction using regular expression for the specified judgment output format. High error rates and low $RS$ are marked red, implying further evaluations (i.e., $RS$, $PC$, and $PF$) are invalid and meaningless because of insufficient accessible data. Therefore, open-source models (e.g., Llama) are not investigated in the paper because of their high error rates under the difficulty and length of our evaluation instances. In contrast, up-to-date commercial models generally demonstrated minimal error and high repetition stability, making them more practical and meaningful for investigations.
Judge MTBench DevBench
@@
Caude-3.5-sonnet 0.96 $\pm$ 0.07 0.82 $\pm$ 0.14 0.01 0.00 0.95 $\pm$ 0.09 0.76 $\pm$ 0.16 0.22 0.00
Claude-3-opus 0.95 $\pm$ 0.08 0.70 $\pm$ 0.19 0.22 0.00 0.96 $\pm$ 0.07 0.69 $\pm$ 0.20 0.29 0.00
Claude-3-sonnet 0.93 $\pm$ 0.11 0.59 $\pm$ 0.22 0.32 0.01 0.95 $\pm$ 0.09 0.71 $\pm$ 0.22 0.23 0.00
Claude-3-haiku 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.18 0.00 0.90 $\pm$ 0.17 0.23 $\pm$ 0.14 0.75 0.00
Gemini-1.5-pro 0.97 $\pm$ 0.09 0.62 $\pm$ 0.19 0.23 0.03 0.87 $\pm$ 0.17 0.84 $\pm$ 0.17 0.03 0.13
Gemini-1.5-flash 1.00 $\pm$ 0.00 0.67 $\pm$ 0.17 0.07 0.00 0.04 $\pm$ 0.08 0.92 $\pm$ 0.39 0.00 0.96
Gemini-1.0-pro 0.89 $\pm$ 0.18 0.57 $\pm$ 0.18 0.30 0.00 0.85 $\pm$ 0.26 0.66 $\pm$ 0.20 -0.05 0.00
o1-mini 0.90 $\pm$ 0.07 0.76 $\pm$ 0.15 -0.04 0.00 0.93 $\pm$ 0.12 0.84 $\pm$ 0.13 -0.07 0.00
GPT-4o 1.00 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.12 0.00 0.98 $\pm$ 0.03 0.80 $\pm$ 0.16 -0.12 0.00
GPT-4-turbo 0.94 $\pm$ 0.10 0.75 $\pm$ 0.16 0.02 0.00 0.97 $\pm$ 0.06 0.79 $\pm$ 0.18 0.16 0.00
GPT-4 0.97 $\pm$ 0.05 0.82 $\pm$ 0.15 0.02 0.00 0.97 $\pm$ 0.05 0.83 $\pm$ 0.15 -0.13 0.00
GPT-3.5-Turbo 0.96 $\pm$ 0.07 0.70 $\pm$ 0.18 0.06 0.00 0.99 $\pm$ 0.02 0.76 $\pm$ 0.18 -0.02 0.00
Llama-2-13B 0.45 $\pm$ 0.22 0.56 $\pm$ 0.33 -0.08 0.55 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-2-7B 0.83 $\pm$ 0.15 0.44 $\pm$ 0.19 0.11 0.17 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 1.00
Llama-3-8B 0.16 $\pm$ 0.18 0.43 $\pm$ 0.39 -0.05 0.84 0.06 $\pm$ 0.09 0.04 $\pm$ 0.09 -0.05 0.94
Figure 3: Baseline comparisons of judges across tasks on MTBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) includes a radar chart comparison by family and an overall baseline bar chart comparison of $PC$. The baseline model is chosen to be gpt-4-0613 for its generally superior performances. Figure (b) demonstrates the $PF$ score of evaluated judges across tasks where the baseline is $PF=0$.
@@
Appendix D Main Findings
Position Bias of Capable Judges are not Mere Random Variations
@@ Appendix D Main Findings
As shown in Table 4, the capable judges on the benchmark tasks, supported by minimal “Error" rates, generally exhibit $RS$ of over 0.85. The up-to-date and most capable models, such as Claude-3.5-Sonnet, Claude-3-Opus, Gpt-4, and GPT-4o, all achieve near-perfect $RS$ of over 0.95 on both benchmarks. These results validate that LLM judgments and the consequently observed position bias are not merely random variations. This increases the confidence that one-time generated judgments by these validated LLMs adequately represent their judging capabilities. Only with this prerequisite, can we ensure further investigations are reasonable and meaningful.
Position Bias Varies by Judge & Task
@@
As shown in Table 2, in general, GPT-4 excels as the best-performing judge model across metrics and benchmarks, hence considered the baseline model in Fig. 3 for $PC$ comparisons. However, certain models achieve comparable or superior performances than GPT-4 on certain tasks. For example, according to Fig. 3, for coding task evaluation, GPT-4o and Claude-3.5-Sonnet are likely more ideal judges. Also, GPT-3.5-Turbo achieves comparable $PC$ as GPT-4, indicating that it may be employed as a cost-effective alternative to coding evaluations.
@@
Similarly, significant variances of $PF$ across tasks are observed. The judges achieving close-to-0 $PF$ in general, such as GPT-4 and Claude-3.5-Sonnet, exhibit varied preference directions across tasks, preferring primacy on some tasks while recency on others. Particularly, o1-mini, while being primacy-preferred on coding, extraction, and math, exhibits almost fair preferences on reasoning, roleplay, and writing tasks. Even for uniformly recency-preferred judges such as Claude-3’s and Gemini-pro’s, the extent of biased preference, as reflected by $PF$ values, varies by task.
@@
Moreover, high position consistency does not guarantee fairness, weighing the trade-offs between position consistency and preference fairness. For example, on coding task evaluations, GPT-4 and GPT-4o achieve the top consistency but are significantly recency-preferred and primacy-preferred, respectively. In comparison, GPT-3.5-Turbo is highly preference fair while having comparable consistency. In this sense, sacrificing consistency for more fairness or vice versa can all be practically useful depending on specific needs. This trade-off between consistency and fairness is particularly crucial when the most consistent model is not the fairest.
Therefore, to sum up, the position bias of LLM judges is task-dependent and there is usually a trade-off between position consistency and preference fairness. In practice, to achieve the most reliable evaluation results, one needs to select task-specific LLMs whose consistency and fairness are in balance. The same discoveries can be concluded from DevBench results, as detailed in Appendix.Sec. K.
Position Bias Correlates to Answer Quality Gap
@@
Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section B.2, we quantify the quality gap ($\delta_{q}$) between answer pairs by the overall win rate ($owr$). Since an absolute tie has 0.5 $owr$, $\delta_{p}$ increases as $owr$ extends from 0.5 to 0 or 1. Fig.4(a) exhibits a significant parabolic shape, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs with larger quality disparities are easier to achieve consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec.J.
@@
Similarly, as shown in Fig. 1(b), judgments generally become more positionally fair as $\delta_{q}$ increases. However, the extent is not as significant as for $PC$. Also, the relationship varies by judge, where some LLMs maintain preference fairness regardless of $\delta_{q}$. For example, as shown in Fig.6, $PF$ of GPT models centered closely around 0 consistently, whereas that of Claude and Gemini-pro models exhibit a conspicuous proportional relationship on MTBench.
@@
These observations align with the right-arrow shape as demonstrated in Fig.4(c), where there is a general trend that judgments become preference fairer as position consistency increases. It also justifies the reasonability of our quantification of positional fairness, as highly position consistent judges are not overly penalized and a perfect $PC$ should result in $PF=0$.
@@
Our investigation also helps to explain the observed verbosity/length bias [11, 17, 3] and self-enhancement bias [12, 3] of LLM judges. From our analysis, LLMs may prefer longer answers (verbosity/length bias) since more capable models usually generate longer solutions with higher qualities. Similarly, in the cases where the same LLM serves as both a candidate and a judge model, judges preferring the answers generated by themselves (self-enhancement bias) are usually the ones who generate responses of higher quality (e.g., GPT-4). In both cases, the observed biased preference may be fundamentally due to a prominent answer quality gap instead of length or self-identity.
Position Bias is Length-Independent
@@
We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the response length of answer-generating models (task output length), and the length of the entire prompt (prompt length). Since the tokenization methods vary by model, we apply the length of the prompt component strings for uniformity and scalability. The analysis from Fig. 4(d)(e)(f) shows that there is no significant and definite relationship between the lengths of prompt components and position bias. Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
Practical Implications for Benchmark Design
@@
We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appedix.Sec. H to gain more insights. Particularly, the disagreement analysis helps to identify the instances where it is easy/difficult for all judges to reach a consensus, reflecting the difficulty of judging them. Based on discoveries of the answer quality gap and LLM agreement/disagreement, our study offers practical implications for evaluator benchmark designs in controlling the difficulty of judging tasks. From our findings, the hard-to-judge instances are the ones where (1) LLMs mostly disagree with each other (2) comparable win rates and minimal quality gaps are observed for the two candidate models (3) prominent position bias exhibits for a majority of judges.
@@
Here is a qualitative analysis of a typical difficult-to-judge example. This instance (MTBench question 120) involved a math problem where both models, vicuna-13b-v1.3 and wizardlm-30b, provided incorrect answers. The problem statement is as follows: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find x such that $f(x)=0$. Methodologically, vicuna-13b-v1.3 erroneously applied the quadratic formula to solve the cubic equation, whereas wizardlm-30b employed an enumeration method but failed to identify the correct answer. This resulted in a 3{A}-4{B}-5{C} choice pattern among the twelve judges, a typical case of significant disagreement.
@@
Qualitatively, We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer. The remaining ones showed preferences either for correctness (which then chose {C} as a tie) or methodological soundness (which then chose {B}, i.e., preferring wizardlm-30b). Notably, in the swapped order setting of the same problem, 8 of the 12 judges encountered position bias. This is a typical illustration of the two responses having a minimal quality gap, leading to difficulty in judging, the likelihood of position bias, and prominent subjectivity during evaluations.
The takeaway message from examples like this is that the number of hard-to-evaluate and highly subjective instances could be reduced for better benchmark designs. Similarly, the trivial-to-judge instances where all judges easily reach a consensus could be reduced among the benchmark population. One could also specify more detailed and definite evaluation criteria to erase the ambiguity of some hard-to-evaluate instances.
Appendix E Conclusion
In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of pairwise comparative LLM-as-a-Judge in terms of repetition stability, position consistency, and preference fairness. We focus on position bias for its prevalent occurrences and significant impact on the reliability of LLM judges’ evaluations. From prior works, the pairwise comparative assessment offers the most consistent and reliable judgments compared to other types of LLM-as-a-Judge and hence becomes our concentration for practicality. The study starts by investigating the repetition stability of LLM judges to validate that the judgments and consequently observed position bias are not mere random variations. Then, in addition to the traditional measurement of positional consistency, we propose a systematic way to quantify the preference fairness of LLM judges. Our comprehensive evaluation of twelve judge models on two benchmarks across 22 tasks reveals significant variance of position bias by judge and task. This stresses the trade-offs between position consistency, preference fairness, and cost-effective practicality when choosing the optimal judges for specific tasks.
The study on the Judge-level, Candidate-level, and Task-level impactful factors on position bias further enhances the understanding of this bias. For example, we are the first the make a direct connection between answer quality gap and position bias, where our discovery reveals that the quality gap between candidate responses to the task question significantly influences the positional consistency and preference fairness of LLM judges. However, position bias is length-independent, meaning that the lengths of prompt components have minimal impact on position bias. In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs. To summarize, insights from this study enhance the understanding of position bias and lay the groundwork for more robust, unbiased evaluation systems.
Appendix F Contribution & Limitation